OpenStudy (anonymous):

Simplify and write the trigonometric expression in terms of sine and cosine: [2+tan^2x/sec^2x]−1=(f(x))^2

OpenStudy (loser66):

break 2 =1+1, then 1+tan^2 = sec^2 after manipulating, you get cos^2 = f(x)^2 hence, f(x) = cos

OpenStudy (anonymous):

thank you!!